OpenStudy (anonymous):

HELP!! FAN & MEDAL, I'll BE SO GRATEFUL!!! Classify each quadratic by matching. 81x^2-16 8a^3-27b^3 4a^2+100 x^2-10xy+25y^2 m^3+64 Perfect square Differences of cubes Sum of Cubes Difference of Squares None of the above Which of the following has the greatest x value for the vertex? Explain your choice. f(x)=x^2+4x+4 f(x)=3(x-1)^2+5 f(x)=-2(x+1)^2+3 f(x)=-3x^2+12x+1 Name 3 methods that can be used to solve quadratic equations (I put factoring, quadratic formula, & completing the square)

jhonyy9 (jhonyy9):

1. (9x)^2 - 4^2 = so what is this ?

jhonyy9 (jhonyy9):

@emilybbydoll are you here ?

OpenStudy (anonymous):

@jhonyy9 Im not sure what it is thats why im asking for help

OpenStudy (anonymous):

the first one is just matching like for ex (im not saying its true) 81x^2-16 would be perfect square

OpenStudy (anonymous):

I just need help on the first two & if you could tell me if the last one is correct?

jhonyy9 (jhonyy9):

so the first one i think will be difference of perfect squares - or not ?

jhonyy9 (jhonyy9):

the second differences of cubes

OpenStudy (anonymous):

the first one is like this, you have to match each one to a term 81x^2-16 8a^3-27b^3 4a^2+100 x^2-10xy+25y^2 m^3+64 Perfect square Differences of cubes Sum of Cubes Difference of Squares None of the above It just all looks the same to me

jhonyy9 (jhonyy9):

the last one is sum of cubes

OpenStudy (anonymous):

okay i see

jhonyy9 (jhonyy9):

what is there again ?

OpenStudy (anonymous):

how about the 3rd and 4th one? If its okay can you please explain to me why, it doesn't have to be intricate or anything

OpenStudy (anonymous):

what we have so far 81x^2-16 Perfect square 8a^3-27b^3 Differences of cubes 4a^2+100 ??? x^2-10xy+25y^2 ??? m^3+64 Sum of Cubes

OpenStudy (anonymous):

The two remaining terms are Difference of Squares & None of the above

jhonyy9 (jhonyy9):

3rd is sum of squares (2a)^2 +10^2 4th x^2 -10xy +25y^2 = (x-5y)^2

jhonyy9 (jhonyy9):

4th perfect square

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (anonymous):

help help help

OpenStudy (anonymous):

Thank you so much! but we already used perfect square once, which one would be None of the above?

OpenStudy (anonymous):

@jhonyy9 are you there?

jhonyy9 (jhonyy9):

3rd is sum of squares but this choice not is there so this will be non of above

OpenStudy (anonymous):

okay so the fourth one would be difference squares since its the last one left?

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

no the 4th is a perfect square (x+5y)^2

jhonyy9 (jhonyy9):

difference squares is the first one

OpenStudy (anonymous):

okay thanks for clearing that up (: can you help me with the second one?

jhonyy9 (jhonyy9):

second what ? exercise ?

OpenStudy (anonymous):

Which of the following has the greatest x value for the vertex? Explain your choice. f(x)=x^2+4x+4 f(x)=3(x-1)^2+5 f(x)=-2(x+1)^2+3 f(x)=-3x^2+12x+1

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

do you know what mean greatest vertex ?

OpenStudy (anonymous):

no please explain

jhonyy9 (jhonyy9):

how can you calculi vertex of a quadratic ?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

please explain to me everything it is I need to know, you've been a life saver so far

jhonyy9 (jhonyy9):

open this ,, http://www.virtualnerd.com/tutorials/?id=Alg1_12_01_0003 " - and here will get a very easy understandably example

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

okay, sorry i lost connection. @jhonyy9

OpenStudy (anonymous):

can you check the last one for me? I just want to know if its correct Name 3 methods that can be used to solve quadratic equations (I put factoring, quadratic formula, & completing the square)